### v2

@@ 4 Algorithmic Considerations
By Thm. 3.6, the most direct approach to computing the set of relevant cuts is to compute, for all distinct pairs $s,t\in V$, the set of all minimal $s,t$-cuts. Even though there may be exponentially many relevant cuts, it is possible to encode them in an intermediary data structure that can be constructed in polynomial time. As shown in [11], the set all minimal $a,b$-cuts for all $a,b$, i.e., the set of relevant cuts, can be computed from a Gomory-Hu tree $T^{*}$ and the directed acyclic graphs introduced by Picard and Queyranne [19] for the edges $st$ of $T^{*}$. A PQ-DAG $D_{s,t}$ is a compact representation of all minimal $s,t$-cuts . It is obtained from $G$ by contracting to a single node each connected component of the residual flow network for an arbitrary maximum $s,t$-flow. Each node in $D_{s,t}$ is identified with the subset of vertices in $V$ from which it was formed. The “closed” node sets in the DAG $D_{s,t}$, i.e., those without out-edges, correspond to the minimal $s,t$-cuts. Both the Gomory-Hu tree $T^{*}$ (for recent advances see [1, 2]) and the PQ-DAG $D_{s,t}$ for any pair $s,t$ can be computed in polynomial time. From this combined data structure, $\mathcal{K}$ can be listed with $O(|\mathcal{K}|n)$ effort [11].
This is reminiscent of the polynomially many “prototypes” for relevant cycles described in [27], from which all relevant cycles can be listed. There is an important difference, however: For relevant cycles, there is a polynomial-time algorithm to count their number based on counting a constrained set of shortest paths [27]. In contrast, counting the number of minimal $s,t$-cuts is #P-complete [21], and thus at least as hard as NP-complete decision problems [25].
The enumeration of all relevant cuts can be sped up considerably with the help of observations made by Gusfield and Noar in [11]. First we need some more notation.
For any $u,v\in V$, we say that $v$ is a successor of $u$ in $D_{s,t}$, if either $u$ and $v$ belong to the same node in $D_{s,t}$ or if the node in $D_{s,t}$ that contains $v$ is a successor of the node in $D_{s,t}$ that contains $u$. Let $D_{s,t}(X,Y)$ with $X,Y\subseteq V$ and $X\cap Y=\emptyset$ denote the PQ-DAG resulting from $D_{s,t}$ by contracting the set $X$ with all its successors, and contracting the set $Y$ with all its predecessors.
@@
Any closed set in $D_{s,t}(X,Y)$ that contains $s$ but not $t$ is a closed set in $D_{s,t}$ that contains $\{s\cup X\}$ but not $\{t\cup Y\}$.
@@
By construction, $s$ is contracted into a supernode with all its successors and $t$ is contracted into a super-node with all its predecessors to create $D_{s,t}$. Accordingly, the super-nodes containing $s$ and $t$ are the only vertices with out-degree 0 and in-degree 0, respectively. For any vertex $u\in V$, $s$ is a successor of $u$ and $t$ is a predecessors of $u$, as a consequence of this construction. Contracting $X$ with all its successors therefore includes $s$ but not $t$ as a successor. Analogously, contracting $Y$ with all its predecessors includes $t$ but not $s$ as a predecessors. Therefore, the super-node of out-degree 0 in $D_{s,t}(X,Y)$ represents $\{s\cup X\}$, and the super-node of in-degree 0 represents $\{t\cup Y\}$. It follows that any closed set in $D_{s,t}(X,Y)$ contains $\{s\cup X\}$ thereby $s$ but not $\{t\cup Y\}$ thereby $t$.
@@
[11, Thm. 2.1] For a fixed pair of vertices $u,v\in V$, let $(x_{1},y_{1}),\ldots,(x_{k},y_{k})$ be the $k$ minimum-weight tree edges along the path between $u$ and $v$ in $T^{*}$. A cut $(V,V^{\prime\prime})$ is a minimum $u,v$-cut if and only if it appears (as a closed set) in at least one PQ-DAG $D_{x_{i},y_{i}}(u,v)$ for some $1\leq i\leq k$, where $D_{x_{i},y_{i}}(u,v)$ is the graph resulting from contracting all successors of $u$ and all predecessors of $v$ in $D_{x_{i},y_{i}}$, respectively, into a single node.
@@
A cut $D\in\mathfrak{D}$ is relevant if and only if it is represented by at least one closed vertex set in a PQ-DAG $D_{a,b}$ derived from an edge in the Gomory-Hu tree $T^{*}$, hence $\{a,b\}\in T^{*}$.
@@
$\Rightarrow$” By Thm. 3.6, if a cut is relevant, it is a weight-minimal $u,v$-cut for some pair of distinct vertices $u$ and $v$. By Lemma 4.3 any weight-minimal cut between two vertices $u$ and $v$ is represented by a closed set in exactly one PQ-DAG $D_{a,b}(u,v)$, where $ab$ is weight minimal along the path from $u$ to $v$ in $T^{*}$. By lemma 4.1 any $a$-$b$ cut in $D_{a,b}(u,v)$ is also a cut in $D_{a,b}$.
@@
$\Leftarrow$” Each edge $e=\{a,b\}$ in a Gomory-Hu tree $T^{*}$ is weight minimal for at least the vertex pair $a$ and $b$ on the path consisting of only the edge $ab$ itself. The PQ-DAG $D_{a,b}$ represents all weight minimal cuts between $a$ and $b$ as closed sets. The closed sets are therefore weight minimal cuts for two distinct vertices $s$ and $t$ and hence relevant as a consequence of Thm. 3.6.
@@
Therefore, it sufficient to compute only the PQ-DAGs associated to edges of the Gomory-Hu tree $T^{*}$. As the edges of a Gomory-Hu tree are labeled by the weights of associated minimal cuts, it is trivial to enumerate the relevant cuts in non-decreasing order using this method.
@@
In addition to the approach of [11] and its derivation, several other algorithms for listing $\mathcal{K}$ are feasible. By Lemma 3.4, one could use the bonds of $G$ as a candidate set and extract $\mathcal{K}$ using the canonical greedy algorithm similar to Horton’s approach to computing minimum cycle bases [12]. Very recently, an algorithm for listing all $(s,t)$-bonds with near linear delay [22] has become available. A disadvantage is that the set of bonds must first be generated and then sorted by non-decreasing weight.
@@
A more appealing alternative is to use one of the approaches for ordered generation of cuts. The first method of this type was proposed in [26]. Later an improved version based on a different solutions of maximum flow problems was published [29]. Since these algorithm produce cuts in the order of non-increasing weights, it is not difficulty to whether a cut $(A,\bar{A})$ is a minimum weight cut for any pair $u\in A$ and $v\in A^{\prime}$. This can be done by keeping a matrix $W$, initialized by $W_{st}=\infty$ for all $s,t$, and updated by the following rule: if $W_{u,v}=\infty$, $u\in A$ and $v\in\bar{A}$ then $W_{u,v}\leftarrow|\partial A|$. In particular, it is not necessary to check for linear independence of a cut with all cuts of strictly smaller weight as is the case for minimal cycle bases. Moreover, the generation process can be terminated as soon as (i) minimum weight cuts have been observed for all vertices, i.e., upon generation of $(A,\bar{A})$, no entries $\infty$ were present in $W$, and (ii) the weight of the current cut is strictly larger than the weight of the previous one, which coincides with the largest finite entry in $W$. The enumeration algorithm in [29] has $\tilde{O}(nm)$ delay, which dominates the $O(n^{2})$ effort for “filtering out” cuts that are not minimal $u,v$-cuts for any pair $u,v$.
@@ 5 Computational Experiments
We implemented both methods derived from [11], enumerating closed sets in PQ-DAGs for all pairs of vertices (GUS-P) and only for PQ-DAGs for edges of $T^{*}$ (GUS-T), as well the algorithm for enumerating cuts by non-increasing weights proposed by [29] (YEH). The cut basis for $10,000$ individual chemical graphs from USPTO was computed for method on a consumer PC (Intel i9-9900K). Times were averaged over $5$ iterations and no parallelization was used. Graphs ranged in size between 4 and 198 nodes with a median of 30 at 3 to 211 edges and a median of 33. As such, chemical graphs are considered to be sparse. Unsurprisingly GUS-T always outperformed both GUS-P and YEH, see Table 1. Consistently, but more pronounced for larger graphs, YEH was significantly faster than GUS-P. Although GUS-P was the fastest algorithm for almost 1% of the graphs, manual inspection showed that this observation is most likely an artifact caused by fluctuations in loading times for very small graphs. Overall, YEH was a factor 2.2 slower then GUS-T to create a cut base. However, YEH is overall more versatile when enumerating cuts.
@@ 6 Concluding Remarks
Beyond the interest in relevant cuts implied by the analogy to relevant cycles, they are also of practical interest e.g. for the design of well-behaved crossover operators in graph-based genetic algorithms. Recently, cut-and-join crossover operators for chemical graphs have been investigated and benchmarked in detail for search tasks in “chemical space”, demonstrating their efficacy [7]. It remains an open question, however, whether the restriction of cuts the set relevant cuts remains sufficient to ensure global reachability in the space of chemical graphs in conjunction with degree-preserving or even more restricted join operations.

### v3

@@ 4 Algorithmic Considerations
As mentioned in the preliminaries, the set of all minimal $s,t$-cuts, which we have shown in Thm. 3.6 to coincide with the set of relevant cuts, can be computed from a Gomory-Hu tree $T^{*}$ and the directed acyclic graphs introduced by Picard and Queyranne [22] for the edges $uv$ of $T^{*}$. Both the Gomory-Hu tree $T^{*}$ (for recent advances, see [1, 2]) and the PQ-DAG $Q_{s,t}$ for any pair $s,t$ can be computed in polynomial time.
From this combined data structure, all the connectivity cuts, i.e., cuts with capacity equal to the connectivity value, can be listed with $O(|E||V|)$ effort [12]. This is feasible as there are at most $\binom{|V|}{2}$ such cuts. This is reminiscent of the polynomially many “prototypes” for relevant cycles described in [30], from which all relevant cycles can be listed. There is an important difference, however: For relevant cycles, there is a polynomial-time algorithm to count their number based on counting a constrained set of shortest paths [30]. In contrast, counting the number of minimal $s,t$-cuts is #P-complete [24], and thus at least as hard as NP-complete decision problems [28].
By Thm. 3.6, the most direct approach to computing the set of relevant cuts is to compute, for all distinct pairs $s,t\in V$, the set of all minimal $s,t$-cuts. We shall see, however, that the enumeration of all relevant cuts can be sped up considerably with the help of observations made by Gusfield and Noar in [12]. First we need some more notation.
For any $u,v\in V$, we say that $v$ is a successor of $u$ in $Q_{s,t}$, if either $u$ and $v$ belong to the same node in $Q_{s,t}$ or if the node in $Q_{s,t}$ that contains $v$ is a successor of the node in $Q_{s,t}$ that contains $u$. Let $Q_{s,t}(X,Y)$ with $X,Y\subseteq V$ and $X\cap Y=\emptyset$ denote the PQ-DAG resulting from $Q_{s,t}$ by contracting the set $X$ with all its successors, and contracting the set $Y$ with all its predecessors.
@@
Any closed set in $Q_{s,t}(X,Y)$ that contains $s$ but not $t$ is a closed set in $Q_{s,t}$ that contains $\{s\cup X\}$ but not $\{t\cup Y\}$.
@@
By construction, $s$ is contracted into a super-node with all its successors and $t$ is contracted into a super-node with all its predecessors to create $Q_{s,t}$. Accordingly, the super-nodes containing $s$ and $t$ are the only vertices with out-degree 0 and in-degree 0, respectively. For any vertex $u\in V$, $s$ is a successor of $u$ and $t$ is a predecessors of $u$, as a consequence of this construction. Contracting $X$ with all its successors therefore includes $s$ but not $t$ as a successor. Analogously, contracting $Y$ with all its predecessors includes $t$ but not $s$ as a predecessor. Therefore, the super-node of out-degree 0 in $Q_{s,t}(X,Y)$ represents $\{s\cup X\}$, and the super-node of in-degree 0 represents $\{t\cup Y\}$. It follows that any closed set in $Q_{s,t}(X,Y)$ contains $\{s\cup X\}$ thereby $s$ but not $\{t\cup Y\}$ thereby $t$.
@@
[12, Thm. 2.1] For a fixed pair of vertices $s,t\in V$, let $(u_{1},v_{1}),\ldots,(u_{k},v_{k})$ be the $k$ minimum-weight tree edges along the path between $s$ and $t$ in $T^{*}$. A cut $(V^{\prime},V^{\prime\prime})$ is a minimum $s,t$-cut if and only if it appears (as a closed set) in at least one PQ-DAG $Q_{u_{i},v_{i}}(s,t)$ for some $1\leq i\leq k$, where $Q_{u_{i},v_{i}}(s,t)$ is the graph resulting from contracting all successors of $s$ and all predecessors of $t$ in $Q_{u_{i},v_{i}}$, respectively, into a single node.
@@
A cut $D\in\mathfrak{D}$ is relevant if and only if it is represented by at least one closed vertex set in a PQ-DAG $Q_{u,v}$ derived from an edge in the Gomory-Hu tree $T^{*}$, hence $\{u,v\}\in T^{*}$.
@@
$\Rightarrow$” By Thm. 3.6, if a cut is relevant, it is a weight-minimal $s,t$-cut for some pair of distinct vertices $s$ and $t$. By Lemma 4.3 any weight-minimal cut between two vertices $s$ and $t$ is represented by a closed set in exactly one PQ-DAG $Q_{u,v}(s,t)$, where $uv$ is a weight minimal edge along the path from $s$ to $t$ in $T^{*}$. By Lemma 4.1 any $u,v$-cut in $Q_{u,v}(s,t)$ is also a cut in $Q_{u,v}$.
@@
$\Leftarrow$” Each edge $e=uv$ in a Gomory-Hu tree $T^{*}$ is weight minimal for at least the vertex pair $u$ and $v$ on the path consisting of only the edge $uv$ itself. The PQ-DAG $Q_{u,v}$ represents all weight minimal cuts between $u$ and $v$ as closed sets. The closed sets are therefore weight minimal cuts for two distinct vertices $s$ and $t$ and hence relevant as a consequence of Thm. 3.6.
@@
Therefore, it is sufficient to compute only the PQ-DAGs associated with edges of a fixed Gomory-Hu tree $T^{*}$ of $G$. As the edges of a Gomory-Hu tree are labeled by the weights of associated minimal cuts, it is trivial to enumerate the relevant cuts in non-decreasing order using this method.
@@
In addition to the approach of [12] and its derivation, several other algorithms for listing relevant cuts are feasible. By Lemma 3.4, one could use the bonds of $G$ as a candidate set and extract relevant cuts using the canonical greedy algorithm similar to Horton’s approach to computing minimum cycle bases [14]. Very recently, an algorithm for listing all $(s,t)$-bonds with near linear delay [25] has become available. A disadvantage is that the set of bonds must first be generated and then sorted by non-decreasing weight.
@@
A more appealing alternative is to use one of the approaches for ordered generation of cuts. The first method of this type was proposed in [29]. Later an improved version based on a different solutions of maximum flow problems was published [32]. Since these algorithms produce cuts in the order of non-increasing weights, it is not difficult to determine whether a cut $(V^{\prime},\bar{V^{\prime}})$ is a minimum weight cut for any pair $s\in V$ and $t\in\bar{V}$. This can be done by keeping a matrix $M$, initialized by $M_{st}=\infty$ for all $s,t$, and updated by the following rule: if $M_{u,v}=\infty$, $u\in V$ and $v\in\bar{V^{\prime}}$ then $M_{u,v}\leftarrow|\partial V^{\prime}|$. In particular, because of Thm. 3.6, it is not necessary to check for linear independence of a cut with the set of all cuts of strictly smaller weight as is the case for minimal cycle bases. Moreover, the generation process can be terminated as soon as (i) minimum weight cuts have been observed for all vertices, i.e., upon generation of $(V,\bar{V}^{\prime})$, no entries $\infty$ were present in $M$, and (ii) the weight of the current cut is strictly larger than the weight of the previous one, which coincides with the largest finite entry in $M$. The enumeration algorithm in [32] has $\tilde{O}(|V||E|)$ delay, which dominates the $O(|V|^{2})$ effort for “filtering out” cuts that are not minimal $u,v$-cuts for any pair $u,v$.
@@ 5 Computational Experiments
Figure 2: Algorithmic performance for random graphs, partitioned by edge probability in the Erdős–Rényi model. Running times were tracked and related to the number of nodes (left) and the number of edges (right) of graphs. Lines represent the average running time, while environments show the 95% percentile interval. All plots present logarithmic y-scales. Regardless of edge density, methods show similar pattern.
Figure 3: Algorithmic performance for chemical graphs. Running times were tracked and related to the number of nodes (left) and the number of edges (right) of graphs. For improved resolution, chemical graphs containing 20 or less atoms are plotted independently (above), contrasting the full dataset (below). Lines represent the average running time, while environments show the 95% percentile interval. Please note that only the lower graphs follow a logarithmic y-scale.
@@
We implemented both methods derived from [12], enumerating closed sets in PQ-DAGs for all pairs of vertices (GUS-P) and only for PQ-DAGs for edges of $T^{*}$ (GUS-T), as well as the algorithm for enumerating cuts by non-increasing weights proposed by [32] (YEH). All methods where implemented in Python (v3.11) utilizing the ubiquitous graph library NetworkX (v. 3.3). All benchmarks where run on a consumer PC (Intel i9-9900K) with sufficient memory and without parallelization. Memory usage overall was negligible for all tested instances and is therefore not reported.
Benchmarks were performed on both artificial and realistic data. As generating cuts is of particular interest in the domain of computational chemistry, we used $10,000$ individual chemical graphs from the USPTO-10K dataset [21], a curated subset of reactions reported in the United States Patent and Trademark Office. Graphs ranged in size between 4 and 198 nodes with a median of 30 at 3 to 211 edges and a median of 33. As such, chemical graphs are considered to be sparse. Random graphs were constructed following the classic Erdős–Rényi model [10], generating 1000 valid instances per parameter variant, i.e. per each 6 to 105 nodes (at step size 2 until 40, then at step size 5) and per edge probability 0.25, 0.5, and 0.75. Only connected graphs where considered. The cut basis was computed for each method and dataset, logging individual running times. Times were averaged over $3$ iterations for chemical graphs.
Random graphs and chemical graphs both show a consistent, but markedly different hierarchy between methods. On all random graphs (see Fig. 2), GUS-T consistently outperformed both GUS-P and YEH, while GUS-P consistent outperformed YEH. Margins widen at increasing edge and node counts. Edge probability does not seems to have any influence on this relationship.
In chemical graphs, unsurprisingly, GUS-T also consistently outperformed both GUS-P and YEH for all but very small graphs (see Fig. 3, Tbl. 1). However, consistently, but more pronounced for larger graphs, YEH was significantly faster than GUS-P. Although GUS-P was the fastest algorithm for almost 1% of the graphs, manual inspection showed that this observation is most likely an artifact caused by fluctuations in loading times for very small graphs. Overall, YEH was a factor 2.2 slower then GUS-T to create a cut basis.
@@ 6 Concluding Remarks
With GUS-T, we derived an effective algorithm to compute exactly all relevant cuts of a graph, based on two central observations: relevant cuts coincide with the cuts that are weight-minimal for some pair of vertices, which in turn are fully encoded as the cuts implied by the edges of a Gomory Hu tree. Picard-Queyranne DAGs yield an effective framework towards this goal. GUS-T consistently outperformed other methods in our benchmarks, regardless of graph properties. Nevertheless, it should be acknowledged that YEH is overall more versatile when enumerating cuts.
GUS-T is in particular well suited to enumerate relevant cuts in chemical applications, where cuts relate to chemical reaction mechanism, such as for cut-and-join crossover operators [7]. It remains an open question, however, whether the restriction of cuts to the set of relevant cuts remains sufficient to ensure global reachability in the space of chemical graphs in conjunction with degree-preserving or even more restricted join operations.